OpenStudy (anonymous):

Which is the equation of a line that passes through the point (3, -4) and has a slope of 2 A.y=2x-11 B.y=2x-10 C.y=2x-4 D.y=2x-2

OpenStudy (anonymous):

B

OpenStudy (anonymous):

(y-(-4))=2(x-3) y+4=2x-6 y=2x-10